### v2

# Instance-Adaptive Keypoint Learning with Local-to-Global Geometric Aggregation for Category-Level Object Pose Estimation

Abstract
Category-level object pose estimation aims to predict the 6D pose and size of previously unseen instances from predefined categories, requiring strong generalization across diverse object instances. Although many previous methods attempt to mitigate intra-class variations, they often struggle with instances exhibiting complex geometries or significant deviations from canonical shapes. To address this challenge, we propose INKL-Pose, a novel category-level object pose estimation framework that enables INstance-adaptive Keypoint Learning with local-to-global geometric aggregation. Specifically, our approach first predicts semantically consistent and geometrically informative keypoints through an Instance-Adaptive Keypoint Generator, then refines them with: (1) a Local Keypoint Feature Aggregator capturing fine-grained geometries, and (2) a Global Keypoint Feature Aggregator using bidirectional Mamba for structural consistency. To enable bidirectional modeling in Mamba, we introduce a Feature Sequence Flipping strategy that preserves spatial coherence while constructing backward feature sequence. Additionally, we design a surface loss and a separation loss to enforce uniform coverage and spatial diversity in keypoint distribution. The generated keypoints are finally mapped to a canonical space for regressing the object’s 6D pose and size. Extensive experiments on CAMERA25, REAL275, and HouseCat6D demonstrate that INKL-Pose achieves state-of-the-art performance and significantly outperforms existing methods.
Index Terms:
Object pose estimation, keypoint learning, geometric feature aggregation.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Object pose estimation is a fundamental yet challenging problem in computer vision and robotics, with broad applications in robotic grasping [1], augmented reality [2], 3D scene understanding [3], and autonomous driving [4]. Existing approaches can be broadly classified into instance-level [5, 6, 7, 8, 9] and category-level [10, 11, 12, 13, 14] methods. Instance-level methods rely on specific CAD models to estimate the poses of known objects. However, their heavy dependence on accurate CAD models makes them difficult to generalize to real-world scenarios, where objects are often novel and lack corresponding models. This severely limits their applicability to unseen instances within the same category or to objects with variations in shape and appearance. As a more flexible alternative, category-level object pose estimation aims to predict the pose and size of previously unseen instances from predefined object categories. Despite its potential, this task remains considerably more challenging due to the significant intra-class variations in object shape, texture, and structure.
Figure 1: Visualization of keypoints in NOCS space predicted by AG-Pose [14] and the proposed INKL-Pose on camera instances with complex geometry. Keypoint colors indicate their NOCS errors, which are linearly mapped within the range $[0,0.2]$ from green (error = 0) to red (error = 0.2). AG-Pose tends to produce clustered and outlier-prone keypoints with higher NOCS errors, whereas INKL-Pose generates spatially diverse and well-distributed keypoints that uniformly cover the object surface with significantly lower errors.
Among recent advances, correspondence-based methods [10, 11, 15, 16, 13] have demonstrated promising performance in category-level object pose estimation. These approaches typically rely on establishing dense 3D-3D correspondences between observed point clouds and category-level canonical spaces, such as the Normalized Object Coordinate Space (NOCS)[10]. In NOCS, all instances from the same category are mapped into a unified 3D coordinate frame, and the object pose and size are recovered by estimating a similarity transformation between the observed object and its canonical reconstruction using the Umeyama algorithm [17]. As a result, the quality of the reconstructed NOCS representation directly influences the final pose estimation accuracy.
To improve the quality of NOCS reconstructions, a common approach is to incorporate category-level shape priors. For example, SPD [11] learns object-specific deformations from predefined category templates and establishes dense correspondences between the observed point cloud and its reconstructed canonical shape. This paradigm has inspired a series of follow-up works [15, 16, 18, 19], which typically employ average shape point clouds as priors and design dedicated deformation networks to better adapt to intra-class variations. However, such priors remain static during both training and inference, limiting the model’s adaptability to diverse and unseen geometries. To address this, Query6DoF [20] introduces class-specific learnable query embeddings that are jointly optimized with the pose estimator, allowing the priors to adapt dynamically during training. More recently, shape-prior-free methods [12, 21, 13, 22, 23, 14] have drawn increasing attention due to their enhanced generalization to novel object instances. Among them, AG-Pose [14] introduces a keypoint-based method that predicts representative keypoints for each instance and establishes keypoint-level correspondences, achieving notable improvements against structural discrepancies. Nevertheless, its keypoint features are derived solely from local attention within fixed neighborhoods and simple global average pooling, limiting their capacity to capture fine-grained geometric details and hindering the model’s ability to perceive global structural consistency. Moreover, the keypoint distribution is regulated by heuristic losses, lacking explicit constraints to ensure uniform surface coverage and sufficient spatial separation. As a result, the predicted keypoints (as illustrated in Fig. 1) often cluster in limited regions, fail to accurately capture the object’s surface coverage, and exhibit substantial NOCS errors, particularly on objects with complex geometries.
To address the above challenges, we propose INKL-Pose, a novel category-level object pose estimation method that enables instance-adaptive keypoint learning with local-to-global geometric aggregation. Specifically, we first extract and fuse RGB and point cloud features using a Pointwise Feature Encoder, and then generate a sparse set of semantically consistent and geometrically informative keypoints through an Instance-Adaptive Keypoint Generator. These keypoints are subsequently refined by two complementary modules: (1) a Local Keypoint Feature Aggregator that captures fine-grained geometric details from spatial neighborhoods, and (2) a Global Keypoint Feature Aggregator that utilizes a bidirectional Mamba architecture to model long-range dependencies while maintaining structural consistency across the object. To handle the unordered and non-causal nature of point clouds, we introduce a Feature Sequence Flipping strategy that facilitates bidirectional perception in the feature space while preserving the original spatial structure. Additionally, we design a surface loss and a separation loss to explicitly encourage uniform coverage and spatial diversity among keypoints. The refined keypoints are then mapped to a canonical space to regress the object’s 6D pose and size. As illustrated in Fig. 1, even on complex camera instances, the keypoints predicted by our INKL-Pose still achieve accurate and consistent surface coverage, and exhibit significantly lower NOCS errors compared to those predicted by AG-Pose[14].
Overall, our contributions are summarized as follows:
  • We propose INKL-Pose, a novel keypoint-based framework for category-level 6D object pose estimation that enables instance-adaptive keypoint learning with local-to-global geometric aggregation.
  • We introduce two complementary modules for keypoint feature refinement: a Local Keypoint Feature Aggregator that captures fine-grained geometric details from spatial neighborhoods, and a Global Keypoint Feature Aggregator that leverages bidirectional Mamba with a Feature Sequence Flipping strategy to enable coherent and globally consistent geometric perception.
  • We design a surface loss and a separation loss to encourage keypoints to be uniformly and diversely distributed across the object surface, thereby enabling more effective capture of semantically consistent and geometrically informative features.
  • Extensive experiments on CAMERA25, REAL275, and HouseCat6D demonstrate that INKL-Pose achieves state-of-the-art performance on both synthetic and real-world benchmarks, significantly outperforming prior methods.
II Related Work
In this section, we discuss the literature most relevant to our work from three perspectives: instance-level object pose estimation, category-level object pose estimation, and state space models.
II-A Instance-Level Object Pose Estimation
Instance-level object pose estimation aims to infer the precise pose of an object based on its corresponding 3D CAD model. Existing approaches in this domain can be broadly divided into three categories: direct regression methods [6, 24], direct voting methods [25, 26] and dense 2D-3D correspondence methods [27, 9]. Direct regression methods utilize end-to-end networks to directly predict the object’s rotation and translation parameters from extracted visual features. Although this approach simplifies the pose estimation pipeline, it suffers from reduced accuracy in cluttered or occluded scenes due to the limitations of direct mapping. Direct voting methods, on the other hand, estimate pose candidates and their corresponding confidence scores at the pixel or point level, ultimately selecting the pose with the highest confidence. While these methods are capable of capturing detailed local information and generally yield strong performance, the dense voting process can be computationally intensive and time-consuming. Alternatively, the third category of instance-level methods adopts a correspondence-based paradigm, which first establishes dense 2D–3D correspondences between RGB images and the complete 3D CAD model, and subsequently estimates the object pose using algorithms such as Perspective-n-Point or least-squares optimization. Unlike direct regression or voting-based approaches, this method explicitly decouples feature matching from pose estimation, thereby enabling more geometry-aware inference when accurate correspondences are available. However, the performance of such methods heavily depends on the availability of precise CAD models and robust feature descriptors, both of which are often difficult to obtain in real-world scenarios.
II-B Category-Level Object Pose Estimation
Unlike instance-level approaches that rely on known CAD models, category-level 6D object pose estimation aims to estimate the pose and size of previously unseen instances from predefined object categories. This setting is inherently more general and challenging due to substantial intra-class variations among object instances. To address these challenges, NOCS [10] introduces the Normalized Object Coordinate Space, which maps different instances to a unified canonical frame, thereby mitigating variations in scale and geometry. The Umeyama algorithm [17] is then applied to recover the object pose via similarity transformation. To better handle intra-class variations, SPD [11] proposes a deformation network that reconstructs object geometry by learning instance-specific deformations from category-level shape priors, followed by dense correspondence matching between the reconstructed canonical model and the observed point cloud. This paradigm has inspired a series of follow-up works [18, 19, 15, 16], which further explore the extraction of more robust features and the development of more efficient deformation strategies. However, the static nature of shape priors during both training and inference limits adaptability and hinders generalization to objects with substantial structural variations.
To overcome this limitation, Query6DoF [20] introduces class-specific learnable queries that are jointly optimized with the pose estimation network, enabling the shape priors to adapt dynamically during training. In addition, shape-prior-free methods [12, 21, 13, 22, 23, 14] have attracted increasing attention for their improved generalization to diverse and unseen object geometries. IST-Net [13] adopts an implicit spatial transformation framework that directly learns correspondences between camera-space and world-space features without relying on explicit priors. CLIPose [12] leverages a vision-language model to extract rich semantic features from both text and images, aligning them with point cloud representations via contrastive learning. TG-Pose [21] introduces topological descriptors and a symmetry-aware reconstruction module to refine object geometry and enhance pose stability. SecondPose [23] integrates geometric features with semantic embeddings from DINOv2 [28], supervised under SE(3)-consistent constraints to improve cross-instance generalization. SpherePose [29] employs a shared spherical proxy representation to address semantic inconsistency arising from shape-dependent canonical spaces. CleanPose [30] integrates causal learning with knowledge distillation to address the spurious correlations introduced by confounding factors in pose estimation. More recently, AG-Pose [14] proposes detecting instance-adaptive keypoints and establishing keypoint-level correspondences to enhance robustness against structural variations across object instances, achieving notable performance improvements.
Our work also focuses on establishing keypoint-level correspondences, but surpasses AG-Pose by introducing a novel keypoint detection method that incorporates explicit constraints on keypoint distribution. Furthermore, our method jointly captures fine-grained geometric details and global structural consistency, enabling the generation of semantically consistent and geometrically informative keypoints. This design provides strong geometric supervision and significantly enhances robustness to intra-class shape variations, ultimately leading to more accurate and generalizable pose estimation.
II-C State Space Models
State Space Models (SSMs) [31], originating from the classical Kalman filter, have attracted increasing interest for their ability to model long-range dependencies with linear computational complexity, while also supporting parallel training. Variants such as S4 [31] introduce diagonal state matrices to reduce memory and computational overhead. Building upon this foundation, Mamba [32] further enhances the SSM framework by incorporating a selective mechanism that compresses context information based on input-dependent parameterization, improving both efficiency and performance. Since the original Mamba operates in a causal manner, where each output token can only attend to preceding tokens, subsequent work such as Vision Mamba [33] introduces bidirectional Mamba to enable information flow in both directions and improve global interaction.
Motivated by these advances, recent works [34, 35, 36, 37, 38] have adapted the Mamba architecture for 3D point cloud analysis. For instance, PointMamba [34] transforms unordered point cloud data into ordered 1D sequences using Hilbert and Trans-Hilbert curves, preserving spatial continuity and semantic relevance. PCM [35] further utilizes bidirectional Mamba with a point sequence flipping strategy to allow each point to access features from any other point in the cloud. Inspired by the success of Mamba in long-range sequence modeling, we explore its potential for category-level object pose estimation. In particular, we leverage bidirectional Mamba to model globally consistent geometric structures across varied object instances, and propose a novel bidirectional input sequence generation strategy tailored to the unordered and non-causal nature of point cloud data.
III Methodology
III-A Overview
In this section, we present the detailed architecture of the proposed INKL-Pose. As illustrated in Fig. 1(a), our objective is to jointly estimate an object’s 3D rotation $R\in\mathbb{R}^{3\times 3}$, 3D translation $t\in\mathbb{R}^{3}$ and size $s\in\mathbb{R}^{3}$ from an input RGB image $\mathcal{I}_{o}$ and its corresponding point cloud $\mathcal{P}_{o}$. Specifically, INKL-Pose consists of five key stages. First, the Pointwise Feature Encoder (Section III-B) extracts instance features from both the RGB image and the point cloud, which are subsequently fused to facilitate cross-modal integration. Second, the fused features are processed by the Instance-Adaptive Keypoint Generator (Section III-C) to predict a sparse set of keypoints and their associated features. Third, the Local Keypoint Feature Aggregator (Section III-D) enriches each keypoint with geometrically distinct local features. Fourth, the Global Keypoint Feature Aggregator (Section III-E) captures globally consistent geometric context across object instances. Finally, the Pose and Size Estimator (Section III-F) takes the enhanced keypoints as input and directly regresses the object’s 6D pose and size.
Figure 2: (a) Overall framework of INKL-Pose [^1] This figure was inspired by the drawing style of CleanPose [30]. We gratefully acknowledge the authors for their work.. (b) The Instance-Adaptive Keypoint Generator dynamically generates keypoints along with corresponding features by assigning weights based on the semantic and geometric information extracted by the Pointwise Feature Encoder. (c) The Local Keypoint Feature Aggregator captures fine-grained local structures by aggregating point cloud neighborhood information. (d) The Global Keypoint Feature Aggregator models long-range dependencies using bidirectional Mamba, extracting globally consistent geometric representations across instances.
III-B Pointwise Feature Encoder
Given an RGB image $\mathcal{I}_{o}$ and its corresponding point cloud $\mathcal{P}_{o}$, the Pointwise Feature Encoder (PFE) extracts informative features from both modalities. Specifically, semantic features $\mathcal{F}_{\mathcal{I}}\in\mathbb{R}^{N\times d_{1}}$ are extracted from the RGB image using DINOv2 [28], while geometric features $\mathcal{F}_{\mathcal{P}}\in\mathbb{R}^{N\times d_{2}}$ are obtained from the point cloud via PointNet++ [39]. To emphasize relative geometric relationships and reduce the influence of absolute spatial positions, the point cloud is preprocessed by subtracting the mean of all coordinates. Although this operation benefits relative geometry modeling, it reduces the network’s sensitivity to absolute positional cues, which are essential for accurate pose estimation. To compensate for this loss, a Multi-Layer Perceptron (MLP) layer is employed to encode the original spatial coordinates into a latent positional embedding $\mathcal{E}_{\mathcal{P}}\in\mathbb{R}^{N\times d_{3}}$. Finally, the semantic features $\mathcal{F}_{\mathcal{I}}$, geometric features $\mathcal{F}_{\mathcal{P}}$, and positional embedding $\mathcal{E}_{\mathcal{P}}$ are concatenated and passed through an additional MLP layer for cross-modal feature fusion, yielding the final pointwise feature representation $\mathcal{F}_{o}\in\mathbb{R}^{N\times d}$, which serves as input for subsequent network modules.
III-C Instance-Adaptive Keypoint Generator
Given the challenges of noise and incomplete observations in cluttered environments, we propose an Instance-Adaptive Keypoint Generator (IAKG) that dynamically generates keypoints from the most informative regions of each object, ensuring robustness and adaptability across diverse instances. Rather than relying on fixed strategies like Farthest Point Sampling (FPS), which select keypoints solely based on spatial distances and fail to adapt to the geometric diversity across instances, our method learns to predict instance-specific keypoints in a data-driven and context-aware manner.
Specifically, we formulate keypoint generation as a learnable query-based process that adapts to both semantic and geometric variations, following the paradigm introduced in AG-Pose [14]. In contrast to AG-Pose, our approach introduces two novel loss functions that explicitly encourage the predicted keypoints to be uniformly distributed and spatially diverse across the object surface. As shown in Fig. 1(b), we first initialize a learnable keypoint query embedding $\mathcal{Q}_{kpt}$, which is initially set to zero, providing a neutral starting point that avoids introducing bias toward specific patterns. This initialization facilitates stable learning dynamics and allows the query vectors to adapt purely based on the input data through the attention mechanism. Then, through multi-head cross-attention and multi-head self-attention mechanisms, the keypoint query vectors interact comprehensively with both the input features $\mathcal{F}_{o}$ and themselves, progressively updating to form instance-specific query embeddings. Finally, we compute the cosine similarity between $\mathcal{Q}_{kpt}$ and $\mathcal{F}_{o}$, and apply a softmax function to normalize the results, yielding the keypoint probability distribution matrix $\mathcal{W}$:
$\mathcal{W}=Softmax\left(\frac{\mathcal{Q}_{kpt}\cdot\mathcal{F}_{o}}{\|\mathcal{Q}_{kpt}\|_{2}\cdot\|\mathcal{F}_{o}\|_{2}+\varphi}\right),$ (1)
where $\varphi$ is a small constant for numerical stability. Based on the matrix $\mathcal{W}$, the keypoint coordinates $\mathcal{P}_{kpt}\in\mathbb{R}^{N_{kpt}\times 3}$ and keypoint features $\mathcal{F}_{kpt}\in\mathbb{R}^{N_{kpt}\times d}$ are computed as weighted averages of input coordinates $\mathcal{P}_{o}$ and fused features $\mathcal{F}_{o}$:
$\mathcal{P}_{kpt}=\mathcal{W}\mathcal{P}_{o},\quad\mathcal{F}_{kpt}=\mathcal{W}\mathcal{F}_{o}.$ (2)
To ensure the predicted keypoints are both well-distributed across the object surface and geometrically diverse, we introduce two complementary loss functions. First, the surface loss $\mathcal{L}_{surf}$ encourages uniform surface coverage by minimizing the Chamfer Distance between the predicted keypoints $\mathcal{P}_{kpt}$ and a set of reference points $\mathcal{P}_{fps}$, which are uniformly sampled from the input point cloud using FPS:
[…]
Second, the separation loss $\mathcal{L}_{sep}$ promotes geometric diversity by encouraging each keypoint to be spatially distinct from its local neighbors. Specifically, it maximizes the average pairwise distance between each predicted keypoint and its $M$-nearest neighbors in Euclidean space:
$\mathcal{L}_{sep}=\frac{1}{\max\left(\frac{1}{N_{kpt}}\sum_{i=1}^{N_{kpt}}\frac{1}{M}\sum_{j=1}^{M}\|\mathcal{P}_{kpt}^{i}-\text{NN}_{j}(\mathcal{P}_{kpt}^{i})\|_{2},\epsilon\right)},$ (4)
where $\text{NN}_{j}(\mathcal{P}_{kpt}^{i})$ denotes the $j$th nearest neighbor of the $i$th keypoint, and $M$ is the number of neighbors considered. In our experiments, we set $M=2$, which provides a lightweight yet effective constraint to mitigate keypoint clustering while maintaining flexibility in spatial layout. The denominator is lower-bounded by a small constant $\epsilon$ = 0.01 to ensure numerical stability.
By jointly optimizing these two loss terms, the predicted keypoints are encouraged to be uniformly distributed and spatially diverse, thus establishing a robust foundation for accurate pose estimation. The effectiveness of these loss functions is demonstrated through ablation studies in Section IV-C.
III-D Local Keypoint Feature Aggregator
While the extracted keypoints effectively localize salient regions of an object, their sparse nature may result in the loss of fine-grained geometric details. To address this limitation, we propose a Local Keypoint Feature Aggregator (LKFA) that enriches each keypoint by integrating geometrically distinct local features from its surrounding neighborhood.
As illustrated in Fig. 1(c), we first design a coordinate encoding function that projects the keypoint coordinates into a high-dimensional feature space, producing positional embeddings $\mathcal{E}_{kpt}\in\mathbb{R}^{N_{kpt}\times d}$. These embeddings are then added to the initial keypoint features $\mathcal{F}_{kpt}$ to enhance spatial awareness:
$\mathcal{E}_{kpt}=MLP(\mathcal{P}_{kpt}),\quad\mathcal{F}_{kpt}=\mathcal{F}_{kpt}+\mathcal{E}_{kpt}.$ (5)
Next, for each keypoint, we construct a local neighborhood in both coordinate and feature spaces using the $K$-nearest neighbor ($K\!N\!N$) algorithm. This yields coordinate neighborhoods $\mathcal{P}_{K}\in\mathbb{R}^{N_{kpt}\times K\times 3}$ and feature neighborhoods $\mathcal{F}_{K}\in\mathbb{R}^{N_{kpt}\times K\times d}$, where $K$ is the neighborhood size. We then concatenate the spatial and feature descriptors and apply standard normalization to obtain fused local group features $\mathcal{F}_{K}^{\prime}\in\mathbb{R}^{N_{kpt}\times K\times(d+3)}$, which jointly encode geometric structure and semantic information:
$\mathcal{F}_{K}^{\prime}=Norm(Concat(\mathcal{F}_{K},\mathcal{P}_{K})).$ (6)
To further enhance the representational capacity of each keypoint, we concatenate the normalized neighbor features $\mathcal{F}^{\prime}_{K}$ with the corresponding central keypoint feature $\mathcal{F}_{kpt}$, resulting in expanded representations $\mathcal{F}^{\prime\prime}_{K}\in\mathbb{R}^{N_{kpt}\times K\times(2\times d+3)}$. This operation allows each keypoint to incorporate both its own feature and contextual information from its spatial neighborhood. We then apply an affine transformation to $\mathcal{F}^{\prime\prime}_{K}$, parameterized by learnable weights $\alpha$ and $\beta$, to capture potential rigid transformations in local geometry and adaptively refine the feature distribution. The transformed feature set ^$\hat{\mathcal{F}}_{K}\in\mathbb{R}^{N_{kpt}\times K\times(2\times d+3)}$ is computed as follows:
$\mathcal{F}_{K}^{\prime\prime}=Concat(\mathcal{F}_{K}^{\prime},\mathcal{F}_{kpt}),\quad\hat{\mathcal{F}}_{K}=\mathbf{\alpha}\odot\mathcal{F}_{K}^{\prime\prime}+\mathbf{\beta},$ (7)
where $\odot$ denotes element-wise multiplication.
A max-pooling operation is then applied across each local neighborhood to aggregate the most discriminative geometric and semantic cues into the central keypoint representation. Finally, an MLP layer is used to refine the output and produce the final local feature representation $\mathcal{F}_{local}\in\mathbb{R}^{N_{kpt}\times d}$:
$\mathcal{F}_{local}=MLP(MaxPool(\mathbf{\hat{\mathcal{F}}}_{K})).$ (8)
By aggregating information from spatially coherent local regions, our LKFA effectively captures geometrically distinct local variations, enhancing the descriptive capacity of each keypoint and laying a solid foundation for robust global perception.
III-E Global Keypoint Feature Aggregator
To complement the geometrically distinct local features captured by the LKFA with geometrically consistent global features across different object instances, we introduce a Global Keypoint Feature Aggregator (GKFA) based on SSMs. This design facilitates robust modeling of global shape structure and improves cross-instance generalization at the category level. To clarify the theoretical underpinnings of our GKFA module, we begin by reviewing the core formulation of SSMs, followed by an introduction to the Mamba architecture built upon them, which serves as the foundation of our global feature modeling.
III-E1 Preliminaries
SSMs [31] were originally introduced in control theory for modeling linear time-invariant (LTI) dynamical systems. The core principle is to map the input signal $x(t)$ to an output signal $y(t)$ through a latent state $h(t)$, whose evolution is governed by both the current state and the input. This process is described by the following system of ordinary differential equations:
$h^{\prime}(t)=Ah(t)+Bx(t),\quad y(t)=Ch(t),$ (9)
where $A$ is the state transition matrix, and $B$ and $C$ are the input and output projection matrices, respectively. Here, $t$ denotes continuous time, and the system captures how the internal dynamics evolve smoothly over time.
To apply SSMs to discrete data such as point cloud sequences or tokens in deep learning, the continuous-time formulation must be discretized. Mamba [32] achieves this via the zero-order hold method, which introduces a time step ($\Delta$) to approximate the continuous dynamics. This leads to a transformation of the continuous matrices $A$ and $B$ into their discrete counterparts $\overline{A}$ and $\overline{B}$, as follows:
$\overline{A}=e^{\Delta A},\quad\overline{B}=(\Delta A)^{-1}(e^{\Delta A}-I)\Delta B.$ (10)
Here, $\Delta$ represents the temporal resolution of the sampled sequence (i.e., the interval between adjacent observations in discrete time). With these discretized parameters, the continuous-time system in Eq. (9) can be reformulated into a recurrent state-space model defined over discrete time steps:
$h_{t}=\overline{A}h_{t-1}+\overline{B}x_{t},\quad y_{t}=Ch_{t},$ (11)
where $t$ denotes the discrete time step. This formulation enables efficient sequence-to-sequence mapping from inputs $x_{t}$ to outputs $y_{t}$, but its inherently sequential nature poses challenges for efficient training.
To support parallelized training, the recurrence can be equivalently expressed in convolutional form:
$\overline{K}=(C\overline{B},C\overline{AB},\ldots,C\overline{A}^{L-1}\overline{B}),\quad y=x\ast\overline{K},$ (12)
where $L$ denotes the length of the input sequence $x$, and $\overline{K}$ represents the global convolution kernel composed of weighted combinations of the discretized system parameters. The operator * denotes the one-dimensional convolution, which applies the kernel $\overline{K}$ over the input sequence $x$ to generate the output $y$. This convolutional formulation allows the model to compute the entire output sequence in parallel, significantly improving training efficiency over the recursive formulation.
III-E2 GKFA Module
As illustrated in Fig. 1(d), to capture global semantic consistency and long-range dependencies among keypoints, we integrate a bidirectional variant of Mamba [35] into the GKFA module. This design enables bidirectional information flow across keypoints, facilitating the learning of globally consistent geometric structures. Traditionally, bidirectional modeling is achieved through a strategy we refer to as Point Sequence Flipping (PSF), where the input point sequence is reversed based on spatial indices. However, such input-order-based reversal is not well suited for point cloud data, as the unordered and non-causal nature of point sets may disrupt spatial and semantic coherence, especially when geometrically unrelated points are juxtaposed in the reversed sequence.
To address this limitation, we propose a Feature Sequence Flipping (FSF) strategy that operates along the feature dimension rather than spatial order. Unlike PSF, FSF preserves the original spatial arrangement of keypoints, while enabling bidirectional modeling through the feature space. As illustrated in Fig. 3, we construct a forward feature sequence by directly using the keypoint features $\mathcal{F}_{local}$, and generate a backward feature sequence by reversing the feature channels of $\mathcal{F}_{local}$. These two sequences are then independently fed into two parallel Mamba blocks. The outputs are aggregated via element-wise summation, followed by a linear projection and residual connection, to produce the final globally aggregated keypoint features $\mathcal{F}_{global}\in\mathbb{R}^{N_{kpt}\times d}$:
$\mathcal{F}_{global}=SSM(\mathcal{F}_{local},FSF(\mathcal{F}_{local}))+\mathcal{F}_{local}.$ (13)
Figure 3: Comparison between traditional Point Sequence Flipping (PSF) and the proposed Feature Sequence Flipping (FSF).
To ensure that the aggregated keypoints retain meaningful and expressive geometric structure, we introduce a reconstruction-based supervision mechanism. Specifically, we employ an MLP layer to regress dense positional offsets from the global keypoint features. Given $\mathcal{F}_{global}\in\mathbb{R}^{N_{kpt}\times d}$, we repeat each feature vector $m=\lfloor N_{rec}/N_{kpt}\rfloor$ times to construct an expanded representation $\mathcal{F}_{global}^{\prime}\in\mathbb{R}^{N_{rec}\times d}$, where $N_{rec}$ is the predefined number of reconstruction points (set to 960 in our experiments). The MLP predicts a set of 3D offsets $\mathcal{O}_{\mathcal{P}}\in\mathbb{R}^{N_{rec}\times 3}$:
$\mathcal{O}_{\mathcal{P}}=MLP(\mathcal{F}_{global}^{\prime}).$ (14)
Similarly, the keypoint coordinates $\mathcal{P}_{kpt}\in\mathbb{R}^{N_{kpt}\times 3}$ are repeated $m$ times to obtain $\mathcal{P}_{kpt}^{\prime}\in\mathbb{R}^{N_{rec}\times 3}$. The reconstructed point cloud $\mathcal{P}_{rec}$ is then computed as:
$\mathcal{P}_{rec}=\mathcal{P}_{kpt}^{\prime}+\mathcal{O}_{\mathcal{P}}.$ (15)
To ensure geometric consistency between the reconstructed and observed point clouds, we adopt a Chamfer Distance-based similarity loss $\mathcal{L}_{sim}$, which measures the bidirectional alignment between $\mathcal{P}_{rec}$ and $\mathcal{P}_{o}$:
$\mathcal{L}_{sim}(\mathcal{P}_{o},\mathcal{P}_{rec})=$ $\,\frac{1}{N_{pts}}\sum_{x_{i}\in\mathcal{P}_{o}}\min_{y_{j}\in\mathcal{P}_{rec}}\|x_{i}-y_{j}\|_{2}^{2}+$ (16)
$\frac{1}{N_{rec}}\sum_{y_{j}\in\mathcal{P}_{rec}}\min_{x_{i}\in\mathcal{P}_{o}}\|y_{j}-x_{i}\|_{2}^{2}.$
This reconstruction supervision encourages each keypoint to act as a generator of local geometry, effectively bridging the gap between sparse semantic abstraction and dense geometric representation. In our ablation studies, we compare different settings for the number of keypoints and find that using 96 (with a repetition factor of 10) achieves the best performance under a fixed reconstruction resolution of 960 points.
To further strengthen the representational capacity of keypoints, we stack multiple LKFA and GKFA modules in a cascaded architecture. Each block refines keypoint features by alternately aggregating fine-grained local geometry and long-range global context. This design allows hierarchical learning across different scales, improving the network’s ability to model complex object structures. As demonstrated by our ablation studies (Section IV-C), using $S$ = 12 stacked blocks yields the best performance.
III-F Pose and Size Estimator
After obtaining semantically consistent and geometrically informative keypoints along with their associated features, we develop a Pose and Size Estimator to directly infer the object’s 6D pose and size. Specifically, we first map the predicted keypoint coordinates to the NOCS space using the ground-truth rotation $\textit{R}_{gt}$, translation $\textit{t}_{gt}$, and scale $\textit{s}_{gt}$, thereby obtaining the ground-truth keypoint coordinates in NOCS space $\mathcal{P}_{kpt}^{gt}\in\mathcal{R}^{N_{kpt}\times 3}$. Subsequently, we predict the NOCS coordinates $\mathcal{P}_{kpt}^{nocs}\in\mathcal{R}^{N_{kpt}\times 3}$ via a self-attention mechanism followed by an MLP layer:
$\mathcal{P}_{kpt}^{gt}=\left\{\frac{1}{\|\textit{s}_{gt}\|_{2}}(x_{i}-\textit{t}_{gt})\textit{R}_{gt}\mid x_{i}\in\mathcal{P}_{kpt}\right\},$ (17)
$\mathcal{P}_{kpt}^{nocs}=MLP(Self\text{-}Attn(\mathcal{F}_{kpt})).$ (18)
To supervise the mapping from keypoints to NOCS, we design a mapping loss $\mathcal{L}_{map}$, employing the smooth $\mathcal{L}_{1}$ loss [14] to constrain the distance between the predicted NOCS coordinates and the ground-truth NOCS coordinates.
By leveraging keypoint-level correspondences, we employ three parallel MLPs to predict the rotation $R_{pre}$, translation $t_{pre}$ and size $s_{pre}$, respectively.
$\mathcal{F}_{pose}=Concat(\mathcal{P}_{kpt},\mathcal{P}_{kpt}^{nocs},\mathcal{F}_{kpt}),$ (19)
$R_{pre},t_{pre},s_{pre}=MLP(\mathcal{F}_{pose}),MLP(\mathcal{F}_{pose}),MLP(\mathcal{F}_{pose}),$ (20)
where $R_{pre}$ and $t_{pre}$ represent the predicted 6D object pose, and $s_{pre}$ represents the predicted object size. We simply use the $\mathcal{L}_{2}$ norm loss function to supervise $R_{pre}$, $t_{pre}$ and $s_{pre}$:
$\mathcal{L}_{pose}=\|R_{pre}-R_{gt}\|_{2}+\|t_{pre}-t_{gt}\|_{2}+\|s_{pre}-s_{gt}\|_{2}.$ (21)
III-G Overall Loss Function
In summary, the overall loss function is as below:
$\mathcal{L}_{all}=$ $\omega_{sep}\mathcal{L}_{sep}+\omega_{surf}\mathcal{L}_{surf}+$ (22)
$\omega_{sim}\mathcal{L}_{sim}+\omega_{map}\mathcal{L}_{map}+\omega_{pose}\mathcal{L}_{pose},$
where $\omega_{sep}$, $\omega_{surf}$, $\omega_{sim}$, $\omega_{map}$, and $\omega_{pose}$ are hyperparameters used to balance the contributions of separation loss, surface loss, similarity loss, mapping loss and pose loss.
IV Experiments
IV-A Experimental Setup
IV-A1 Datasets
Following previous works [14, 29], we evaluate the effectiveness of the proposed INKL-Pose on three widely used benchmarks: CAMERA25 [10], REAL275 [10], and the more challenging HouseCat6D [40] dataset. CAMERA25 is a large-scale synthetic dataset covering six object categories, containing 300k synthetic RGB-D images with objects rendered on virtual backgrounds, of which 25k images are used for evaluation. REAL275 serves as a real-world counterpart to CAMERA25, covering the same six categories. The training set comprises 4.3k images captured from 7 scenes, while the test set includes 2.75k images from 6 different scenes, with three distinct object instances per category. HouseCat6D is a more complex, real-world multimodal dataset featuring 10 object categories. The training set contains 20k RGB-D images collected across 34 indoor scenes, and the test set includes 3k images from 5 additional scenes. Notably, this dataset poses greater challenges due to its marker-free capture process, dense clutter, and severe occlusion.
IV-A2 Evaluation Metrics
Following previous works [12, 14, 29], we evaluate the performance of INKL-Pose using two widely adopted metrics.
  • IoUX. This metric calculates the 3D Intersection-over-Union (IoU) between the predicted and ground-truth bounding boxes. A prediction is considered correct if the IoU exceeds a specified threshold. Following standard protocol, we report IoU50 and IoU75 on CAMERA25 and REAL275, and IoU25 and IoU50 on the more challenging HouseCat6D dataset.
  • $n$°$m$cm. This metric evaluates the 6D pose accuracy by jointly considering rotation and translation errors. A prediction is regarded as correct if the rotation error is less than $n$° and the translation error is below $m$ cm. Following standard protocol, we report results under four commonly used thresholds: 5°2cm, 5°5cm, 10°2cm, and 10°5cm across all three datasets.
TABLE I: Quantitative comparisons With state-of-the-art methods on the REAL275 dataset[10]. The best results are in bold.
  Method Prior REAL275 ($\uparrow$)
$\mathrm{IoU_{50}}$ $\mathrm{IoU_{75}}$ $5^{\circ}2\mathrm{cm}$ $5^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$
SPD [ECCV’20][11] 77.3 53.2 19.3 21.4 43.2 54.1
SGPA [ICCV’21][15] 80.1 61.9 35.9 39.6 61.3 70.7
GPV-Pose [CVPR’22][41] 83.0 64.4 32.0 42.9 55.0 73.3
DPDN [ECCV’22][16] 83.4 76.0 46.0 50.7 70.4 78.4
GPT-COPE [TCSVT’23][18] 82.0 70.4 45.9 53.8 63.1 77.7
HS-Pose [CVPR’23][42] 82.1 74.7 46.5 55.2 68.6 82.7
IST-Net [ICCV’23][13] 82.5 76.6 47.5 53.4 72.1 80.5
Query6DoF [ICCV’23][20] 82.5 76.1 49.0 58.9 68.7 83.0
VI-Net [ICCV’23][22] - - 50.0 57.6 70.8 82.1
CatFormer [AAAI’24][19] 83.1 73.8 47.7 53.7 69.0 79.5
CLIPose [TCSVT’24][12] - - 48.7 58.3 70.4 85.2
TG-Pose [TMM’24][21] 82.6 76.2 49.8 59.0 71.7 86.6
SecondPose [CVPR’24][23] - - 56.2 63.6 74.7 86.0
AG-Pose [CVPR’24][14] 83.8 77.6 56.2 62.3 73.4 81.2
AG-Pose (DINOv2) [CVPR’24][14] 84.1 80.1 57.0 64.6 75.1 84.7
SpherePose [ICLR’25][29] 84.0 79.0 58.2 67.4 76.2 88.2
Ours 84.0 80.6 63.2 68.3 80.9 88.0
TABLE II: Quantitative comparisons with state-of-the-art methods on the CAMERA25 dataset[10]. The best results are in bold.
  Method Prior CAMERA25 ($\uparrow$)
$\mathrm{IoU_{50}}$ $\mathrm{IoU_{75}}$ $5^{\circ}2\mathrm{cm}$ $5^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$
SPD [ECCV’20] [11] 93.2 83.1 54.3 59.0 73.3 81.5
SGPA [ICCV’21] [15] 93.2 88.1 70.7 74.5 82.7 88.4
GPV-Pose [CVPR’22] [41] 92.9 86.6 67.4 76.2 - 87.4
GPT-COPE [TCSVT’23] [18] 92.5 86.9 70.4 76.5 81.3 88.7
HS-Pose [CVPR’23] [42] 93.3 89.4 73.3 80.5 80.4 89.4
Query6DoF [ICCV’23] [20] 91.9 88.1 78.0 83.1 83.9 90.0
VI-Net [ICCV’23] [22] - - 74.1 81.4 79.3 87.3
TG-Pose [TMM’24] [21] 93.4 89.7 74.5 81.9 81.1 90.4
CatFormer [AAAI’24] [19] 93.5 89.9 74.9 79.8 85.3 90.2
CLIPose [TCSVT’24] [12] - - 74.8 82.2 82.0 91.2
AG-Pose [CVPR’24] [14] 94.1 91.7 77.1 82.0 85.5 91.6
AG-Pose (DINOv2) [CVPR’24] [14] 94.2 92.5 79.5 83.7 87.1 92.6
SpherePose [ICLR’25] [29] 94.8 92.4 78.3 84.3 84.8 92.3
Ours 94.4 92.9 81.0 85.0 88.1 93.4
Figure 4: Qualitative comparisons between INKL-Pose with AG-Pose[14] on the REAL275 dataset[10], where green and red represent the predicted and ground-truth bounding boxes, respectively. Our INKL-Pose demonstrates superior precision in pose estimation.
IV-A3 Implementation Details
We utilize the off-the-shelf Mask R-CNN [43] to generate instance masks and resize all input images to 224$\times$224 pixels. For each corresponding object instance, we uniformly sample 1,024 points from the segmented point cloud. To enhance robustness, we apply data augmentation strategies consistent with prior works [14, 16], including random rotations (within 0–20 degrees), translations (in the range of –0.02 to 0.02 meters), and scalings (from 0.8 to 1.2). For network configuration, the feature dimensions are set as follows: $d_{1}$ = 128, $d_{2}$ = 128 , $d_{3}$ = 64 , and the fused feature dimension is $d=256$. In the local feature aggregation stage, the neighborhood size $K$ is set to 4. Additionally, the entire feature aggregation module consists of $S$ = 12 stacked stages, each comprising both local and global aggregation components to progressively refine the keypoint features. In the keypoint generation stage, the network predicts 96 keypoints per object instance. For keypoint supervision, we sample $N_{fps}$ = 120 reference points to ensure sufficient surface coverage and spatial alignment with the predicted keypoints. In the separation loss, the number of nearest neighbors is set to $M$ = 2, which provides a lightweight yet effective constraint that mitigates keypoint clustering while preserving flexibility in spatial distribution. In the dense point cloud reconstruction stage, each keypoint is repeated 10 times, yielding $N_{rec}$ = 960 reconstructed points. This setting closely matches the resolution of the input point cloud (1,024 points), providing a stable foundation for reconstruction supervision without incurring excessive computational cost.
Regarding optimization, the weights of the loss terms are set as {$w_{sep},w_{surf},w_{sim},w_{map},w_{pose}$} = {10.0, 10.0, 15.0, 2.0, 0.3}. The model is trained using the Adam optimizer [44] with a cyclical learning rate schedule [45], where the learning rate varies between $2\times 10^{-5}$ and $5\times 10^{-4}$. All experiments are conducted on a single NVIDIA RTX 4090D GPU with a batch size of 64. For CAMERA25 and REAL275, we train the model using both synthetic and real-world data with a synthetic-to-real ratio of 3:1. For HouseCat6D, the model is trained exclusively on its corresponding dataset.
IV-B Comparison with State-of-the-Art Methods
IV-B1 Results on REAL275
Table I reports the performance of INKL-Pose compared with state-of-the-art methods on the REAL275 dataset [10]. As shown in the table, our method exhibits excellent performance on most metrics. Specifically, INKL-Pose achieves 84.0% on IoU50, 80.6% on IoU75, 63.2% on 5°2cm, 68.3% on 5°5cm, 80.9% on 10°2cm, and 88.0% on 10°5cm, demonstrating its strong capability in both pose and size estimation under real-world conditions. Notably, INKL-Pose operates directly on RGB-D input and does not rely on any pre-defined shape priors. Compared with CatFormer [19], the state-of-the-art prior-based method, INKL-Pose achieves significant improvements 15.5% and 14.6%, under the strict 5°2cm and 5°5cm metrics, respectively. Furthermore, compared to SpherePose [29], the current best-performing shape-prior-free method, INKL-Pose yields performance gains of 5.0% and 0.9% under the strict 5°2cm and 5°5cm metrics, respectively. In addition, compared to our strong baseline AG-Pose [14], INKL-Pose achieves notable improvements of 6.2% and 3.7% on 5°2cm and 5°5cm metrics, respectively, highlighting the effectiveness of our fine-to-global geometric aggregation strategy. Qualitative comparisons in Fig. 4 further validate the superiority of INKL-Pose.
IV-B2 Results on CAMERA25
Table II reports the performance of INKL-Pose compared with state-of-the-art methods on the CAMERA25 dataset. As shown in the table, INKL-Pose achieves outstanding accuracy across nearly all evaluation metrics. Although its IoU50 metric reaches 94.4%, slightly lower than SpherePose [29] (achieving 94.8%), it significantly outperforms all existing approaches on the remaining metrics, especially under the more rigorous pose accuracy thresholds. Specifically, compared to prior-based methods, INKL-Pose maintains a clear advantage without relying on pre-defined shape priors. In addition, when compared with AG-Pose [14], the current state-of-the-art shape-prior-free method, INKL-Pose achieves notable gains of 1.5% and 1.3% under the strict 5°2cm and 5°5cm metrics, respectively. These results, along with the strong performance observed on the REAL275 dataset, provide compelling evidence of the robustness and effectiveness of our proposed local-to-global geometric feature aggregation framework.
IV-B3 Results on HouseCat6D
To further validate the generalization capability of the proposed INKL-Pose, we conduct experiments on the more challenging real-world dataset HouseCat6D [40]. As reported in Table III, INKL-Pose achieves state-of-the-art performance across all evaluation metrics. Compared with the current state-of-the-art AG-Pose [14], our method surpasses it under all evaluation metric, with gains of 3.9% on 5°2cm, 5.3% on 5°5cm, 6.2% on 10°2cm, and 7.6% on 10°5cm. Moreover, Fig. 5 presents a detailed per-category comparison of 3D IoU, rotation, and translation accuracy between INKL-Pose and AG-Pose. The results reveal that INKL-Pose outperforms AG-Pose on the majority of metrics, particularly excelling in rotation estimation. Furthermore, Fig. 6 presents qualitative results of the predicted poses, further showcasing the strong generalization ability of INKL-Pose in complex real-world scenarios.
Additionally, Fig. 2 offers intuitive visual comparisons between INKL-Pose and other state-of-the-art methods on both the REAL275 and HouseCat6D datasets. These results collectively demonstrate the superior generalization capability and practical applicability of our method in real-world scenarios.
TABLE III: Quantitative comparisons with state-of-the-art methods on the HouseCat6D Dataset[40]. The best results are in bold.
  Method HouseCat6D ($\uparrow$)
IoU25 IoU50 $5^{\circ}2\mathrm{cm}$ $5^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$
GPV-Pose[41] 74.9 50.7 3.5 4.6 17.8 22.7
VI-Net[22] 80.7 56.4 8.4 10.3 20.5 29.1
SecondPose[23] 83.7 66.1 11.0 13.4 25.3 35.7
AG-Pose[14] 82.4 66.0 11.5 12.6 37.4 42.5
AG-Pose (DINOv2)[14] 88.1 76.9 21.3 22.1 51.3 54.3
SpherePose[29] 88.8 72.2 19.3 25.9 40.9 55.3
Ours 90.8 81.9 25.2 27.4 57.5 61.9
Figure 5: Per-category comparison between INKL-Pose and AG-Pose[14] on the HouseCat6D dataset[40], including mean Average Precision (mAP) for 3D IoU, rotation, and translation.
Figure 6: Visualization results of INKL-Pose on the HouseCat6D dataset [40]. Green bounding boxes denote ground truth, while red bounding boxes represent the predictions from our INKL-Pose. The predicted boxes show high consistency with the ground truth.
IV-C Ablation Studies
In this section, we conduct comprehensive ablation studies to evaluate the contributions of key components in the proposed INKL-Pose. All experiments are performed on the REAL275 [10] dataset.
Figure 7: Performance comparison between INKL-Pose and state-of-the-art methods on the REAL275 and HouseCat6D real-world datasets [^2] This figure was inspired by the drawing style of CleanPose [30]. We gratefully acknowledge the authors for their work..
IV-C1 Effect of the Number of Keypoints
We begin by evaluating how the number of predicted keypoints influences the overall pose estimation performance. As shown in Table IV, even with only 16 sparse keypoints, INKL-Pose achieves 56.0% accuracy under the 5°2cm metric, which is already comparable to many existing state-of-the-art methods. As the number of keypoints increases, the performance steadily improves due to a stronger capacity for capturing fine-grained geometric details. The best result is achieved with 96 keypoints, reaching 63.2% on 5°2cm, which is adopted as the default setting in our framework.
IV-C2 Effect of Keypoint Supervision
In the Instance-Adaptive Keypoint Generator (IAKG) module (Section III-C), we introduce two auxiliary losses: $\mathcal{L}_{surf}$ and $\mathcal{L}_{sep}$, to promote uniform surface coverage and enforce spatial diversity among keypoints. To validate the effectiveness of this loss design, we conduct two types of experiments: (1) an ablation study by selectively removing each loss term, and (2) a comparative study by replacing our losses with those used in AG-Pose [14]. As shown in Table IV, removing $\mathcal{L}_{surf}$ results in a 2.8% drop under the 5°2cm metric, demonstrating its role in promoting even keypoint distribution and suppressing outliers. In contrast, removing $\mathcal{L}_{sep}$ leads to a larger performance degradation of 16.3%, highlighting its importance in preventing keypoint clustering. When both losses are removed, the performance drops sharply by 44.7%, confirming the necessity of the proposed supervision strategy. In the comparative setting, we replace our loss functions with those adopted in AG-Pose, which include an object-centric distribution loss and a global diversity term. Compared to these alternatives, our proposed losses provide stronger and more stable geometric supervision. Specifically, $\mathcal{L}_{surf}$ leverages bidirectional Chamfer Distance to ensure comprehensive surface coverage, while AG-Pose employs a one-sided nearest-distance constraint. Meanwhile, $\mathcal{L}_{sep}$ introduces localized repulsion based on nearest neighbors, enabling finer-grained control over spatial dispersion than the global pairwise formulation used in AG-Pose. Together, these results demonstrate the superiority of our loss design in guiding keypoint learning with both semantic consistency and geometric structure.
IV-C3 Effect of Positional Embeddings
We further analyze the contribution of positional information in enhancing feature representation. In INKL-Pose, positional encodings are introduced at two distinct stages: the embedding $\mathcal{E}_{\mathcal{P}}$ in the Pointwise Feature Encoder (Section III-B), which compensates for lost absolute position cues, and $\mathcal{E}_{kpt}$ in the LKFA module (Section III-D), which improves local spatial awareness. As reported in Table IV, removing either embedding leads to a significant performance drop, highlighting their complementary roles in facilitating accurate geometric perception.
IV-C4 Effect of the LKFA and GKFA Modules
We evaluate the individual and combined contributions of the LKFA and GKFA modules introduced in Section III-D and Section III-E, respectively. As shown in Table IV, removing the LKFA causes a 2.1% drop in the 5°2cm metric, primarily due to the loss of fine-grained geometric features. In contrast, removing the GKFA results in a larger decline of 3.0%, indicating its role in maintaining global consistency. When both modules are removed, performance drops by 3.7%, demonstrating their complementary importance in accurate pose estimation.
IV-C5 Effect of the Number of LKFA and GKFA Modules
We investigate the impact of the number of stacked keypoint feature aggregation blocks, where each block comprises both a Local Keypoint Feature Aggregator (LKFA) and a Global Keypoint Feature Aggregator (GKFA). As shown in Table IV, increasing the number of blocks from $S=4$ to $S=12$ consistently improves performance across all evaluation metrics. These results demonstrate that deeper cascaded aggregation enhances both local and global keypoint representations by progressively refining fine-grained geometric details and capturing long-range structural dependencies. However, further increasing the number of blocks to $S=16$ results in performance degradation, as excessive stacking introduces feature redundancy and oversmoothing, which reduce keypoint distinctiveness and impair pose estimation accuracy. Overall, $S=12$ offers an optimal balance between accuracy and efficiency, effectively capturing both local geometry and global structure without incurring unnecessary complexity.
IV-C6 Effect of Neighborhood Size
To evaluate the influence of neighborhood size $K$ in the LKFA module, we vary the number of nearest neighbors used during local feature aggregation. As shown in Table IV, when $K=1$, the model only uses the center point feature, which limits its ability to capture surrounding geometric and semantic context. Increasing $K$ enhances local structure modeling, with optimal performance achieved at $K=4$. Further increasing $K$ introduces redundant information, leading to no additional gains.
IV-C7 Effect of Bidirectional Mamba and Feature Sequence Flipping
We also examine the contributions of the bidirectional Mamba architecture and the proposed Feature Sequence Flipping (FSF) strategy within the GKFA module (Section III-E), with results summarized in Table IV. As the results show, replacing the bidirectional Mamba with its unidirectional counterpart (Uni-Mamba) leads to consistent performance degradation, confirming the effectiveness of bidirectional modeling for global feature aggregation. Furthermore, we compare FSF with the commonly used Point Sequence Flipping (PSF) strategy. PSF amplifies the non-causal nature of point clouds by reversing the sequence along spatial indices, thereby disrupting the alignment between feature order and geometric structure. This misalignment leads to a clear performance drop, even falling below the unidirectional baseline. In contrast, FSF retains the original point order and achieves bidirectional information flow in the feature space, significantly improving robustness to the unordered and non-causal nature of point cloud data.
TABLE IV: Ablation studies on the REAL275 dataset[10]. Default configurations are colored in gray, with best results highlighted in bold.
  $N_{kpt}$ $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
16 56.0 64.3 75.3 85.5
32 57.6 66.0 76.1 87.1
64 59.6 66.7 78.4 87.4
96 63.2 68.3 80.9 88.0
128 59.3 67.7 77.3 87.2
(a) Effect of the number of keypoints.
  Method $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
Full ($\mathcal{L}_{sep}$+$\mathcal{L}_{surf}$) 63.2 68.3 80.9 88.0
w/ $\mathcal{L}_{sep}$ 60.8 67.7 77.8 86.2
w/ $\mathcal{L}_{surf}$ 46.9 53.0 69.1 78.8
w/ AG-Pose losses 60.9 67.0 79.9 87.2
None 18.5 20.4 43.7 53.3
(b) Effect of keypoint supervision.
  Method $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
Full 63.2 68.3 80.9 88.0
w/o $\mathcal{E}_{\mathcal{P}}$ 60.4 67.1 77.6 86.2
w/o $\mathcal{E}_{kpt}$ 59.1 66.6 76.9 85.8
(c) Effect of positional embeddings.
  Method $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
Full 63.2 68.3 80.9 88.0
w/o LKFA 61.1 67.2 80.5 88.4
w/o GKFA 60.2 67.6 77.1 86.6
None 59.5 67.7 78.0 88.3
(d) Effect of the LKFA and GKFA modules.
  $S$ $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
4 59.6 67.5 76.9 86.3
8 61.4 67.0 79.0 86.8
12 63.2 68.3 80.9 88.0
16 60.7 66.9 79.6 87.6
(e) Effect of the number of LKFA and GKFA modules.
  $K$ $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
1 59.6 67.5 77.6 87.0
2 61.1 68.0 79.2 88.0
4 63.2 68.3 80.9 88.0
6 59.7 68.1 77.9 88.4
8 59.9 67.9 77.0 87.1
(f) Effect of neighborhood size.
  Method $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
Uni-Mamba 60.0 66.6 78.5 86.8
PSF 58.2 65.2 77.0 85.2
FSF 63.2 68.3 80.9 88.0
(g) Effect of bidirectional Mamba and feature sequence flipping.
V Conclusion
In this paper, we propose INKL-Pose, a novel keypoint-based framework for category-level object pose estimation that enhances generalization to unseen object instances through instance-adaptive keypoint learning and local-to-global geometric aggregation. Specifically, we introduce an Instance-Adaptive Keypoint Generator that predicts semantically consistent and geometrically informative keypoints, along with two complementary modules for feature refinement: a Local Keypoint Feature Aggregator that captures fine-grained geometric details and a Global Keypoint Feature Aggregator based on bidirectional Mamba with a Feature Sequence Flipping strategy for structure-aware global perception. Furthermore, we design a surface loss and a separation loss to explicitly supervise keypoint distribution, providing strong geometric guidance for robust pose estimation. Extensive experiments on CAMERA25, REAL275, and HouseCat6D demonstrate that INKL-Pose achieves state-of-the-art performance across both synthetic and real-world datasets, significantly outperforming existing approaches. In future work, we will focus on extending our framework to dynamic and temporal settings, enabling sequential pose tracking, and explore uncertainty-aware keypoint modeling to further improve robustness under occlusion and sensor noise.
Acknowledgments
This work was supported by the Wuhan Young Talent Program (Wuhan Talent Initiative) (Grant No. X24D090), the Hubei Science and Technology Innovation Team (Grant No. T2023009), the Innovation Fund of the Hubei Key Laboratory of Intelligent Robots, Wuhan Institute of Technology (Grant No. HBIRL202409), and the Scientific Research Fund of Wuhan Institute of Technology (Grant No. 24QD07).